f(x)=1/(4x+3) Domain
Answer by nerdybill(7384)

f(x)=1/(4x+3) Domain
.
Domain specifies all possible values of x:
The only value you must exclude is when the denominator is zero:
4x+3 = 0
4x = -3
x = -3/4
x = -0.75
.
In set-notation:
(-oo, -0.75) U (-.75, +oo)
where
oo stands for infinity